OpenStudy (3psilon):

Vertical asymptote of f(x) = tan2x ? How would you solve to determine it?

OpenStudy (anonymous):

Hint: Use the definition of \(\tan(x)\), and see where \(\cos(2x)=0\).

OpenStudy (3psilon):

sin/cos?

OpenStudy (anonymous):

Eeeyup.

OpenStudy (3psilon):

@LolWolf I'm getting like .78 . Do i just solve by Taking the inverse cos of 0 and dividing by 2?

OpenStudy (3psilon):

I'm a little rusty on this

OpenStudy (anonymous):

Nah, try to find it analytically, use the unit circle to get a general result.

OpenStudy (3psilon):

Think you can help me out with that? lol

OpenStudy (3psilon):

@LolWolf

OpenStudy (anonymous):

Sure. All right, using this: http://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/300px-Unit_circle_angles_color.svg.png We find that \[ \cos(x)=0 \]When: \[ x=\frac{\pi}{2} \]We also know that, every interval of \(\pi\), the value will repeat (since it is zero for every half-circle), thus, our final solution is: \[ \cos(x)=0, x=\pi k+\frac{\pi}{2} \]Where, \(k\in \mathbb{Z}\). Using this, we have that, vertical asymptotes are present in \(\tan(x)\) when \(x\) is the above value.

OpenStudy (3psilon):

wait so it isn't sin(2x)/cos(2x)??

OpenStudy (3psilon):

I do get the above answer though. My precal knowledge is coming back :) But what happened to the 2x?

OpenStudy (anonymous):

Oh, crap, thanks, yes, I forgot about that, haha. So, for that, then: \[ \cos(2x)=0\implies\\ \cos(y)=0, y=2x\implies x=\frac{y}{2} \]With this, we simply take the expression above and divide it by two, to find: \[ x=\frac{2\pi k+\pi}{4}, k\in \mathbb{Z} \]

OpenStudy (anonymous):

So, there we go, that's for \(\tan(2x)\).